ABC Company has set the following standards in order to produce one unit of its single product: standard quantity standard price direct materials 2.7 yards ?? per yard direct labor 2 hours $14 per hour variable overhead 2 hours ?? per hour During August, ABC Company spent $133,770 to purchase direct materials and had direct labor totaling $189,440. During August, ABC Company used 16,000 yards of direct materials in the production of units. ABC had 2,500 yards of materials in its August 1 direct material inventory and had 1,200 yards yards remaining in its direct materials inventory at August 31. ABC Company reported the following variances for August: Direct material price variance .............. $12,495 unfavorable Direct labor rate variance .................. $24,240 unfavorable Total direct labor variance ................. $6,560 favorable Variable overhead spending variance ......... $1,900 unfavorable Variable overhead efficiency variance ....... $20,900 favorable Calculate ABC's direct material quantity variance for August. If the variance is favorable, place a minus sign in front of your answer (i.e., -5000). If the variance is unfavorable, simply enter your answer as a number (i.e., 5000).
Answer #1

Solution:

Actual quantity of material purchased = Material used + Ending inventory - Beginning inventory

= 16000 + 1200 - 2500 = 14700 yards

Actual price of material = $133,770 / 14700 = $9.10 per yard

Direct material price variance = $12,495 Unfavorable

(SP - AP) * AQ Purchased = -$12,495

(SP - $9.10) * 14700 = -$12,495

Standard price of material = $8.25 per yard

Nos of units produced = 7000 units

Standard quantity for actual production = 7000*2.70 = 18900 yard

Direct material quantity variance = (SQ - AQ) * SP = (18900 - 16000) * $8.25 = $23,925 Unfavorable